Timetable Thursday, September 12

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304

09.00 – 09.45 Plenary Talk (Chair: Josef Stoer) Jean-Charles Gilbert: OPINeL, an interior-point code for nonlinear optimization

09.50 – 10.20 (Chair: Johannes Jahn) Vladimir F. Demyanov: Mathematical diagnostics problems via nonsmooth analysis (Chair: Zdzislaw Naniewicz) Andrzej Myśliński: Primal-dual approach for shape optimization of viscoelastic dynamic contact problems

10.20 – 10.50 Ewa Gora: The penalty functions and Lagrange principle for optimality. Singular extremal problems Irena Pawłow: Control problem for a nonlinear thermoelasticity system

10.50 – 11.15  Coffee break

11.15 – 12.00 Plenary talk (Chair: Andreas Griewank) Krzysztof C. Kiwiel: Applications oriented developments in subgradient and bundle methods for nondifferentiable optimization

12.05 – 12.35 (Chair: Elza Farkhi) Robert Baier: Directed sets as a tool for the visualization of set-valued derivatives (Chair: Ivan Ginchev) Alberto Zaffaroni: Degrees of efficiency and degrees of minimality

12.35 – 13.05 Marcin Studniarski: Necessary and sufficient conditions for weak sharp minima of nonsmooth nonconvex functions Andreas Hamel: Minimal set theorems and applications

13.05 – 14.30  Lunch

14.30 - 15.15 Plenary talk (Chair: Stefan Rolewicz) Andrzej Ruszczyński: Stochastic dominance and mean-risk models

Timetable Thursday, September 12 (continued)

Time  Room LG 1, HS 1  Room LG 1, HS 2  Room LG 1, HS 304
15.15 – 15.45  Coffee break

15.45 – 16.15 Minisymposium (Chair: Sabine Pickenhain) Martin Gugat: Optimal control of flow in networks governed by PDE’s (Chair: Ivo Nowak) Andrzej Mackiewicz: Nonmonotone tensor method strategies for difficult systems of nonlinear equations (Chair: Nikolai Osmolovskii) Natalia Balashevich: Real-time optimization of control systems

16.15 – 16.45 Oskar von Stryk: New frontiers of optimal control for hybrid dynamical systems Christian Zillober: On the impact of large scale mathematical programming on the modeling of topology optimization problems Shamsha Aipanov: Optimal control problems for dynamic systems with fixed endpoints of the trajectories

16.45 – 17.15 Frédéric Bonnans: Bilevel optimization for optimal control problems involving design parameters Andreas Griewank: (to be announced)

Anita Kripfganz: Methods of optimal control theory in plane convex geometry

18.30  Conference Dinner